# Workload Estimation for Unknown Tasks: A Survey of Machine Learning Under Distribution Shift

### v1

@@ IV-C Test-time Adaptation
(Test-time Adaptation [109]) Given a source domain $D_{S}$ and a learning task $L_{S}$, as well as a target domain $D_{T}$ and learning task $L_{T}$, test-time adaptation aims to improve the learning of a target predictive function $f_{T}(x)$ based on an existing predictive function $f_{S}(x)$, where $D_{S}\neq D_{T}$.
@@
Test-time adaptation techniques seek to overcome domain shifts by separating the training process into two distinct phases: (1) training on $D_{S}$ data, and (2) updating that model with $D_{T}$ data. This separation enables the resulting model to adapt at test-time by allowing for the continuous updating of model parameters, as shown in Figure 1 [109].
@@
Prior work applied test-time adaptation techniques to egocentric task recognition [110] and adaptive policy optimization for assistive robots [111]. However, test-time adaptation techniques have not been widely applied to HRI application domains. Nevertheless, these techniques hold promise because they overcome the limitation of training on source and target domain data simultaneously. These techniques can be categorized into as either self-supervised, or entropy-based.
@@
Self-supervised techniques use a dual-output neural network and a two-phase training procedure [109]. Labeled data from the $D_{S}$ undergoes some deterministic transformation (e.g., image rotation). The neural network is trained on this transformed data to predict both the original label (e.g., image class) and the corresponding transformation (e.g., rotation angle). The unlabelled data from $D_{T}$ undergoes the same deterministic transformation. The shared feature extraction model’s parameters are updated based on the output of the self-supervised proxy learning tasks. This two-phase procedure decouples training from adaptation and enables the model to adapt at test-time. Many techniques take inspiration from this work (e.g., [112, 113, 114]); however, the development of an application domain relevant proxy task is non-trivial, limiting the generalizability of self-supervised techniques.
@@
Entropy-based techniques eliminate the need for a self-supervised proxy task by using a batch processing adaptation phase (e.g., [115, 116, 117]). Batches of data from $D_{T}$ are passed through the shared feature extractor model, and the Shannon Entropy [118] is calculated for each batch. The adaptation phase updates the shared feature extractor to minimize the average entropy over all batches [119].
@@
Fundamentally, test-time adaptation is domain adaptation without access to target domain data during training. Therefore, test-time adaptation exhibits one-to-one portability, high model complexity, and low adaptability. It is difficult to definitively state whether these techniques are applicable to workload estimation for unknown tasks, as the relationship between model complexity and adaptability for these techniques is unknown. Prior work has primarily focused on deep learning architectures, but these techniques are general enough to apply to shallow neural networks. It is possible that smaller networks may possess sufficient adaptability to account for real world variability, but the efficacy of test-time adaptation on these networks has not been thoroughly evaluated.
IV-D Continual Learning
@@
(Continual learning [120]) Given a sequence of domains $D$ = $D_{s_{1}},...,D_{s_{n}}$, where $n>0$, and a sequence of learning tasks $T$ = $L_{s_{1}},…,L_{s_{n}}$, where learning task $L_{s_{i}}$ corresponds to domain $D_{s_{i}}$, continual learning aims to optimize the predictive function $f_{i}(x)$ using knowledge from previously seen domains $D_{j}\in D$, where $i\leq j$.
@@
These techniques gradually acquire knowledge to optimize performance on the latest learning task, and some techniques add the constraint of avoiding performance degradation on all prior tasks [121]. Continual learning is also referred to as lifelong learning, sequential learning, or online learning [122]. There is a wide range of techniques, including knowledge distillation (e.g., [123, 124, 125]), memory-based methods (e.g., [126, 127]), and parameter isolation (e.g., [128, 129]).
Continual learning techniques are one of the more popular techniques within HRI, having been successfully applied to problems of human task recognition using wearable sensors (e.g., [130, 131, 132, 133]), affective computing (e.g., [8, 9, 134]), and adaptive human-robot interactions (e.g., [8, 135, 136]).
@@
Conventional continual learning assumes the new learning tasks arrive one at a time, and the data in the domains is stationary [137]. Thus, continual learning techniques can be trained offline as the number of tasks is known a priori. These techniques are referred to as multi-head techniques, as separate models are trained for each learning task [138]. However, knowing all the tasks ahead of time is not always realistic. Other continual learning techniques make no assumptions about the distribution of learning tasks, and seek to identify the task online [138]. These techniques are referred to as single-head techniques, since a single model is trained for all tasks.
@@
Learning tasks within a sequence can either be independent or dependent. Independent learning tasks are separate and may not share similarities. Dependent learning tasks represent a sub-sequence of learning tasks that must be executed in a particular order, where learning task $L_{i}$ holds key information about learning task $L_{i+1}$. Many of these techniques rely heavily on deep neural networks, similar to both domain adaptation and transfer learning techniques [139, 140].
@@
The model complexity of continual learning techniques varies; however, there is evidence that model complexity is not directly tied to performance, as the machine learning model choice is problem specific (e.g., Gaussian processes [136], shallow neural networks [127], and deep networks [122]). Many techniques exhibit low adaptability, requiring thousands of examples to perform well [122]. Portability is challenging to define for continual learning, as it exhibits many-to-many portability, but does so only in a sequential fashion. Many-to-many portability is desirable, but the sequential nature is an unnecessary constraint. Additionally, some continual learning methods only optimize performance on the latest learning tasks, while other optimize performance on all prior learning tasks. Portability and adaptability are key qualities required to estimating workload successfully for unknown tasks; thus, continual learning is a subpar option.
Figure 2: Comparison of applying the technique Model Agnostic Meta-learning [141] to: (a) Domain Generalization, sub-problems experience a domain shift, and (b) Few-shot Learning, sub-problems experience a domain and label shift.
IV-E Domain Generalization
@@
Domain generalization considers the setting where training data from multiple source domains is characterized by the same features that are sampled from different distributions [142]. Specifically, domain generalization is the process of transferring knowledge acquired from several related source domains and applying it to previously unknown domains [143], as indicated in Figure 1.
Definition IV.5
@@
(Domain Generalization [142]) Given a set of source domains $DS$ = $D_{s_{1}},…,D_{s_{n}}$ where $n>0$, a target domain, $D_{t}$, a set of source tasks $LS$ = $L_{s_{1}},…,L_{s_{n}}$ where $L_{s_{i}}\in LS$ corresponds with $D_{s_{i}}\in DS$, and a target task $L_{t}$ that corresponds to $D_{t}$, domain generalization aims to optimizes a target predictive function $f_{t}(x)$ in $D_{t}$, where $D_{t}\notin DS$.
@@
The goal is to leverage unique cross-domain information to improve the machine learning model’s generalization. These methods assume the construction of a joint distribution of source domains aggregates away key information, and that the varying quality and relevance of data across source domains helps inform a model in unknown target domains. A key feature of domain generalization is the lack of labeled target domain samples. Domain generalization techniques can be broadly classified into three categories: domain-invariant representation learning (e.g., [144, 145, 146]), data manipulation (e.g., [147, 148, 149]), and learning strategy (e.g., [150, 151, 152]).
@@
The primary technique for domain-invariant representation learning is domain alignment, which seeks to minimize the difference across source domains, such that a latent representation consisting of only essential features is learned [153]. Domain alignment techniques measure this difference in many ways, including minimizing moments [154], contrastive loss [155], KL divergence [156], or maximum mean discrepancy [157]. Prior work combined domain alignment metrics with domain-adversarial learning [157]. The primary difference between domain-adversarial techniques for domain generalization and domain adaptation is that the source/target discriminator must differentiate between multiple source domains.
@@
Data manipulation techniques seek to artificially enhance the size and diversity of a dataset via data augmentation or generation. Data augmentation performs deterministic transformations over existing data to create new examples [158]. However, data augmentation is not domain-agnostic, as useful transformations are specific to an application domain. Data generation uses generative machine learning models (e.g., Variational Autoencoders [147], Generative Adversarial Networks [148]) to create completely new training examples. These techniques are domain-agnostic, but the underlying models are hard to train and require a high training data volume [159]. These techniques cannot be applied to workload estimation, as generating realistic physiological metrics that correctly correspond to the workload values is non-trivial.
@@
Learning strategy-based domain generalization techniques seek to improve cross domain generalization through novel training methodologies, including: 1) Self-supervsied learning 2) Ensemble learning, and 3) Meta-learning. Self-supervised learning techniques for domain generalization do not differ substantially from self-supervised techniques for test-time adaptation (see Section III.C). Ensemble learning is the process of training multiple machine learning models with varying weight initializations or training data splits, then aggregating the models’ outputs to make predictions. These extensively studied techniques are notoriously computationally expensive [160] and are not viable solutions for real-time HRI domains.
@@
Meta-learning trains a machine learning model on data across source domains, such that that model can be efficiently updated to operate in arbitrary target domains [141]. Broadly, meta-learning is ‘learning to learn’ [161]. Domain generalization typically leverages parameter-based meta-learning techniques (e.g., Model Agnostic Meta-Learning (MAML) [141]) that seeks to construct a neural network that can be adapted based on a few examples (see Figure 2).
@@
A meta-model, parameterized by $\theta$, is trained to predict an initialization for the parameters of another model, $\phi_{i}$, that performs the learning task $L_{i}$. Both $\theta$ and $\phi_{i}$ typically represent a neural network’s weights, but can also represent a subset of the network’s weights, or the another model’s parameters. Each source domain $D_{i}$ and learning task $L_{i}$ is deconstructed into a collection of mini-datasets. These mini-datasets are defined by the support set, $S_{i}$, and the query set, $Q_{i}$, which are small training and test sets, respectively. The support set consists of a few (e.g., 1, 5, 10) examples that the meta-model uses to predict the labels for the query set. The task-specific model’s error is twice differentiated and backpropagated through both models to update $\theta$. This problem formulation constructs the learning process, such that the model learns how to adapt its parameters based on a few examples. This technique is popular in meta-learning [161], and is the foundation for many techniques (e.g, [162, 163, 164, 165, 166]). An alternative approach learns the parameters of an optimizer. An LSTM was trained to learn the exact optimization process of a neural network, such that the new network converged using only a few samples [167]. The technique Meta-Stochastic Gradient Descent combined the LSTM approach and MAML to learn an initialization of the parameters, and an optimal optimizer [168].
@@
The majority of domain generalization techniques exhibit many-to-many portability by construction. Data manipulation is characterized by high model complexity, due to its reliance on generative deep learning architectures [169]. Additionally, these techniques exhibit a range of adaptability, where some instances require several thousands of examples for adaptation [170], and others only require a few examples [171]. MAML and other meta-learning techniques exhibit high adaptability and are general enough to apply to shallow neural networks; thus, exhibiting a low model complexity. Therefore, meta-learning based domain generalization techniques may be applicable to workload estimation for unknown tasks; however, these techniques have been shown to be computationally expensive at both training and inference times [162].
IV-F Few-shot Learning

### v2

@@ 4.3 Test-time Adaptation
(Test-time Adaptation) Given a source domain $D_{S}$ and a learning task $L_{S}$, as well as a target domain $D_{T}$ and learning task $L_{T}$, test-time adaptation aims to improve the learning of a target predictive function $f_{T}(x)$ based on an existing predictive function $f_{S}(x)$, where $D_{S}\neq D_{T}$ (Sun et al., 2020).
@@
Test-time adaptation techniques seek to overcome domain shifts by separating the training process into two distinct phases: (1) training on $D_{S}$ data, and (2) updating that model with $D_{T}$ data. This separation enables the resulting model to adapt at test-time by allowing for the continuous updating of model parameters, as shown in Figure 1 (Sun et al., 2020).
@@
Prior work applied test-time adaptation techniques to egocentric task recognition (Plananamente et al., 2022) and adaptive policy optimization for assistive robots (He et al., 2023). However, test-time adaptation techniques have not been widely applied to HRI application domains. Nevertheless, these techniques hold promise because they overcome the limitation of training on source and target domain data simultaneously. These techniques can be categorized into as either self-supervised, or entropy-based. Enumerating the algorithmic details for each of these techniques will highlight how they may be used in future human workload estimation algorithms.
@@
Self-supervised techniques use a dual-output neural network and a two-phase training procedure (Sun et al., 2020). Labeled data from the $D_{S}$ undergoes some deterministic transformation (e.g., image rotation). The neural network is trained on this transformed data to predict both the original label (e.g., image class) and the corresponding transformation (e.g., rotation angle). The unlabelled data from $D_{T}$ undergoes the same deterministic transformation. The shared feature extraction model’s parameters are updated based on the output of the self-supervised proxy learning tasks. This two-phase procedure decouples training from adaptation and enables the model to adapt at test-time. Many techniques take inspiration from this work (e.g., Azimi et al., 2022; Chen et al., 2023; Segu et al., 2023); however, the development of an application domain relevant proxy task is non-trivial, limiting the generalizability of self-supervised techniques.
@@
Entropy-based techniques eliminate the need for a self-supervised proxy task by using a batch processing adaptation phase (e.g., Liu et al., 2021; Zhang et al., 2022, 2023a). Batches of data from $D_{T}$ are passed through the shared feature extractor model, and the Shannon Entropy (Shannon, 1948) is calculated for each batch. The adaptation phase updates the shared feature extractor to minimize the mean entropy over all batches (Wang et al., 2020a). Other techniques use the same approach, but minimize Shannon Entropy from $D_{S}$ data and $D_{T}$ data (Jung et al., 2023; Seto et al., 2024). Prior work demonstrated that minimizing entropy in this way allows the feature extraction model to learn domain-invariant features to improve generalization to new learning tasks (Adachi et al., 2023). Recent work has extended entropy-based techniques to incorporate varying methods for comparing data points between $D_{S}$ and $D_{T}$ (e.g., Song et al., 2023; Niu et al., 2022; Zhang et al., 2023b).
@@
Fundamentally, test-time adaptation is domain adaptation without access to target domain data during training. Therefore, test-time adaptation exhibits one-to-one portability, high model complexity, and low adaptability. It is difficult to definitively state whether these techniques are applicable to workload estimation for unknown tasks, as the relationship between model complexity and adaptability for these techniques is unknown. Prior work has primarily focused on deep learning architectures, but these techniques are general enough to apply to shallow neural networks. It is possible that smaller networks may possess sufficient adaptability to account for real-world variability, but the efficacy of test-time adaptation on these networks has not been evaluated.
4.4 Continual Learning
@@
(Continual learning) Given a sequence of domains $D$ = $D_{s_{1}},\dots,D_{s_{n}}$, where $n>0$, and a sequence of learning tasks $T$ = $L_{s_{1}},\dots,L_{s_{n}}$, where learning task $L_{s_{i}}$ corresponds to domain $D_{s_{i}}$, continual learning aims to optimize the predictive function $f_{i}(x)$ using knowledge from previously seen domains $D_{j}\in D$, where $i\leq j$ (Chen and Liu, 2018).
@@
These techniques gradually acquire knowledge to optimize performance on the latest learning task, and some techniques add the constraint of avoiding performance degradation on all prior tasks (Smith et al., 2023). Continual learning is also referred to as lifelong learning, sequential learning, or online learning (Mai et al., 2022). There is a wide range of techniques, including knowledge distillation (e.g., Chen et al., 2022a; Wu et al., 2019; Zhu et al., 2021), memory-based methods (e.g., Cai et al., 2021; Derakhshani et al., 2021), and parameter isolation (e.g., Lomonaco et al., 2021; Serra et al., 2018).
Continual learning techniques are one of the more popular techniques within HRI, having been successfully applied to problems of human task recognition using wearable sensors (e.g., Hasan and Roy-Chowdhury, 2015; Ye et al., 2019; Ashry et al., 2020; Leite and Xiao, 2022), affective computing (e.g., Churamani et al., 2020; Gao et al., 2023; Churamani et al., 2022)), and adaptive HRI (e.g., Churamani et al., 2020; Lesort et al., 2020; Spaulding et al., 2021). Further, probabilistic continual learning techniques are common in the cognitive science and psychology communities to construct human decision making models (e.g., Austerweil and Griffiths, 2013; Collins and Frank, 2013; Liew et al., 2016; Kanwal et al., 2022; Navarro et al., 2006; Niv, 2019). Specifically, Dirichlet process mixture models that function as a non-parametric Bayesian clustering algorithm that can recognize existing classes and detect out-of-distribution classes (Li et al., 2019b). Recent work demonstrated that these techniques can be combined with deep-learning models to improve performance for computer vision tasks (e.g., Jerfel et al., 2019; Xu et al., 2024; Willes et al., 2022).
@@
Conventional continual learning assumes the new learning tasks arrive one at a time, and the data in the domains is stationary (De Lange et al., 2021). Thus, continual learning techniques can be trained offline, as the number of tasks is known a priori. These techniques are called multi-head techniques, as separate models are trained for each learning task (Chaudhry et al., 2018). However, knowing all the tasks a priori is not always realistic. Other continual learning techniques make no assumptions about the distribution of learning tasks, and seek to identify the task online (Chaudhry et al., 2018). These single-head techniques train a single model is trained for all tasks.
@@
Learning tasks within a sequence can either be independent or dependent. Independent learning tasks are separate and may not share similarities. Dependent learning tasks represent a sub-sequence of learning tasks that must be executed in a particular order, where learning task $L_{i}$ holds key information about learning task $L_{i+1}$. Many of these techniques rely heavily on deep neural networks, similar to both domain adaptation and transfer learning techniques (Mundt et al., 2023; Ao and Fayek, 2023).
@@
The model complexity of continual learning techniques varies; however, there is evidence that model complexity is not directly tied to performance, as the machine learning model choice is problem specific (e.g., Gaussian processes (Spaulding et al., 2021), shallow neural networks (Derakhshani et al., 2021), and deep networks (Mai et al., 2022)). Many techniques exhibit low adaptability, requiring thousands of examples to perform well (Mai et al., 2022). Portability is challenging to define for continual learning, as it exhibits many-to-many portability, but does so only in a sequential fashion. Many-to-many portability is desirable, but the sequential nature is an unnecessary constraint. Additionally, some continual learning methods only optimize performance on the latest learning tasks, while other optimize performance on all prior learning tasks. Portability and adaptability are key qualities required to estimating workload successfully for unknown tasks; thus, continual learning is a subpar option.
4.5 Domain Generalization
@@
Domain generalization considers the setting where training data from multiple source domains is characterized by the same features, but sampled from different distributions (Li et al., 2018a). Specifically, domain generalization is the process of transferring knowledge from several related source domains and applying it to previously unknown domains (Muandet et al., 2013), shown in Figure 1.
Definition:.
@@
(Domain Generalization) Given a set of source domains $DS$ = $D_{s_{1}},\dots,D_{s_{n}}$ where $n>0$, a target domain, $D_{t}$, a set of source tasks $LS$ = $L_{s_{1}},\dots,L_{s_{n}}$ where $L_{s_{i}}\in LS$ corresponds with $D_{s_{i}}\in DS$, and a target task $L_{t}$ that corresponds to $D_{t}$, domain generalization aims to optimizes a target predictive function $f_{t}(x)$ in $D_{t}$, where $D_{t}\notin DS$ (Li et al., 2018a).
@@
The goal is to leverage unique cross-domain information to improve the machine learning model’s generalization. These methods assume that constructing a joint distribution of source domains aggregates away key information, and that the varying quality and relevance of data across source domains helps inform a model in unknown target domains. A key feature of domain generalization is the lack of labeled target domain samples. This setting exhibits natural similarities to human workload estimation for unknown tasks, as novel tasks will incur a novel physiological response not encountered in the training dataset (i.e., domain shift). Additionally, it is unlikely that labeled data will be available when attempting to estimate workload for unknown tasks in real-world environments. Domain generalization techniques can be broadly classified into three categories: domain-invariant representation learning (e.g., Suh et al., 2023; Li et al., 2022a; Gu et al., 2023b), data manipulation (e.g., Ilse et al., 2020; Li et al., 2021a; Yue et al., 2019), and learning strategy (e.g., Khandelwal and Yushkevich, 2020; Sicilia et al., 2023; Liu et al., 2020).
@@
The primary technique for domain-invariant representation learning is domain alignment, which seeks to minimize the difference across source domains, such that a latent representation consisting of only essential features is learned (Hu et al., 2020). Domain alignment techniques measure this difference in many ways, including minimizing moments (Ghifary et al., 2016), contrastive loss (Kim et al., 2021), or maximum mean discrepancy (Li et al., 2018b). A key difference between domain-adversarial techniques for domain generalization and domain adaptation is that the discriminator must differentiate between multiple source domains.
@@
Data manipulation techniques seek to artificially enhance the size and diversity of a dataset via data augmentation or generation. Data augmentation performs deterministic transformations over existing data to create new examples (Otálora et al., 2019). However, data augmentation is not domain-agnostic, as useful transformations are specific to an application domain. Data generation uses generative machine learning models (e.g., Variational Autoencoders (Ilse et al., 2020), Generative Adversarial Networks (Li et al., 2021a)) to create completely new training examples. These techniques are domain-agnostic, but the underlying models are hard to train and require a high data volumes (Goodfellow et al., 2020). These techniques cannot be applied to workload estimation, as generating realistic physiological metrics that correctly correspond to the workload values is non-trivial.
@@
Learning strategy-based domain generalization techniques seek to improve generalization through novel training methodologies, including: 1) Self-supervised learning 2) Ensemble learning, and 3) Meta-learning. Self-supervised learning techniques for domain generalization do not differ substantially from self-supervised techniques for test-time adaptation. Ensemble learning is the process of training multiple machine learning models with varying weight initializations or training data splits, then aggregating the models’ outputs to make predictions. These techniques are notoriously computationally expensive (Zhou, 2012) and are not viable solutions for real-time HRI domains.
@@
Meta-learning trains a machine learning model on data across source domains, such that that model can be efficiently updated to operate in arbitrary target domains (Finn et al., 2017). Broadly, meta-learning is ‘learning to learn’ (Hospedales et al., 2021). Domain generalization typically leverages parameter-based meta-learning techniques (e.g., Model Agnostic Meta-Learning (MAML) (Finn et al., 2017)) that seeks to construct a neural network that can be adapted based on a few examples (see Figure 2).
@@
A meta-model, parameterized by $\theta$, is trained to predict an initialization for the parameters of another model, $\phi_{i}$, that performs the learning task $L_{i}$. Both $\theta$ and $\phi_{i}$ typically represent a neural network’s weights, but can also represent a subset of the network’s weights, or the another model’s parameters. Each source domain $D_{i}$ and learning task $L_{i}$ is deconstructed into a collection of mini-datasets. These mini-datasets are defined by the support set, $S_{i}$, and the query set, $Q_{i}$, which are small training and test sets, respectively. The support set consists of a few (e.g., 1, 5, 10) examples that the meta-model uses to predict the labels for the query set. The task-specific model’s error is twice differentiated and backpropagated through both models to update $\theta$. This problem formulation constructs the learning process, such that the model learns how to adapt its parameters based on a few examples. This technique is popular in meta-learning (Hospedales et al., 2021), and is the foundation for many techniques (e.g., Antoniou et al., 2018; Bertinetto et al., 2018; Finn et al., 2018; Nichol et al., 2018; Yoon et al., 2018). An alternative approach learns the parameters of an optimizer. An LSTM was trained to learn the exact optimization process of a neural network, such that the new network converged using only a few samples (Ravi and Larochelle, 2016). The technique Meta-Stochastic Gradient Descent combined the LSTM approach and MAML to learn an initialization of the parameters, and an optimal optimizer (Li et al., 2017).
@@
The majority of domain generalization techniques exhibit many-to-many portability by construction. Data manipulation is characterized by high model complexity, due to its reliance on generative deep learning architectures (Zhou et al., 2020). Additionally, these techniques exhibit a range of adaptability, where some instances require several thousands of examples for adaptation (Xu et al., 2021), and others only require a few examples (Chen et al., 2022b). MAML and other meta-learning techniques exhibit high adaptability and are general enough to apply to shallow neural networks; thus, exhibiting a low model complexity. Therefore, meta-learning based domain generalization techniques may be applicable to workload estimation for unknown tasks; however, these techniques have been shown to be computationally expensive at both training and inference times (Antoniou et al., 2018).
Figure 2: A human-robot team conducting two tasks, where each task is characterized by a unique distribution across the seven workload components. These tasks were implemented as a part of an experimental design for developing a multi-dimensional workload estimation method for unknown tasks (Bhagat Smith et al., 2024).
4.6 Few-shot Learning